# The Fourier spectral approach to the spatial discretization of quasilinear hyperbolic systems

### v1

@@ 1 Introduction
Assuming that the system (1.1) has only polynomial nonlinearities may seem an over-restrictive assumption. This assumption is motivated by two considerations. Firstly, the Fourier spectral method may be efficiently implemented only within this framework; see Remark 1.2 below. Secondly, as far as we know, the literature lacks a theory analogous to para-differential calculus (see e.g.  [19, Chapter 5]) on Sobolev spaces of periodic functions, which would provide the composition estimates in Proposition A.4 for general composition functions.
We shall assume further additional structural assumptions on the system (1.1), depending on the needs. In particular we shall always consider Friedrichs-symmetrizable systems, which guarantees the hyperbolicity of the system, and local-in-time well-posedness of the initial-value problem in $L^{2}$-based Sobolev spaces of sufficiently high regularity index, $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ with $s>d/2+1$ (see Subsection 1.2). For simplicity, we consider in this work $2\pi$-periodic functions in all spatial directions. The results extend straightforwardly to more general (periodic) frameworks, and analogous results in the $n$-dimensional Euclidean space could be obtained with some simple adaptations.
@@
Let ${\sf P}_{N}\colon L^{2}((2\pi\mathbb{T})^{d})^{n}\to\mathcal{T}_{N}^{n}$ be the $L^{2}$-projection operator onto $\mathcal{T}_{N}^{n}$: ${\sf P}_{N}=\operatorname{Diag}(P_{N}(D))$, where ${\sf P}_{N}$ is a Fourier multiplier with symbol $P_{N}=\bm{1}_{\llbracket-N,N\rrbracket^{d}}$. We refer to ${\sf P}_{N}$ as a sharp low-pass filter. For any $0\leq r\leq s$ and $\bm{U}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ Sobolev space of order $s$, ${\sf P}_{N}$ satisfies (see e.g. [5, (5.1.10),(5.8.4)])
@@
We do not discuss in this work the full time-space discretization of (1.1), that is well-suited numerical time integrators for (1.3). The implementation we used for numerical experiments is described in more detail in Subsection 4.2.
@@
Our results will depend on the structure of the system. As we shall see, the solution to (1.3) converges towards the corresponding solution to the problem (1.1) (assuming sufficient regularity) whenever the system is symmetric. If the system is only symmetrizable, the situation is more complicated. In order to deal with this situation we consider smooth low-pass filters, ${\sf S}_{N}\colon L^{2}((2\pi\mathbb{T})^{d})^{n}\to\mathcal{T}_{N}^{n}$ where ${{\sf S}_{N}=\operatorname{Diag}(S_{N}(D))}$ is a Fourier multiplier with symbol $S_{N}(\cdot)=S(\cdot/N)$ where $S\in{\mathcal{C}}^{1}(\mathbb{R}^{d})$ is even and satisfies
$\begin{cases}S(\bm{k})=1&\text{if $\max_{j=1,\ldots,d}|k_{j}|\leq 1/2$,}\\ S(\bm{k})=0&\text{if $\min_{j=1,\ldots,d}|k_{j}|\geq 1$,}\\ S(\bm{k})\in{[}0,1{]}&\text{otherwise.}\end{cases}$
and $\left|\cdot\right|\nabla S^{1/2}\in L^{\infty}(\mathbb{R}^{d})$. When $d=1$, an example of such a function could be $S_{1}(\cdot)\coloneq\max\big{(}0,\min\big{(}1,2-2|\cdot|\big{)}\big{)}^{2}$. When $d\geq 2$, one can set $S_{d}(k_{1},\dots,k_{d})\coloneq S_{1}(k_{1})\times\cdots\times S_{1}(k_{d})$. The advantage of such smooth low-pass filters is that —contrarily to the sharp low-pass filter— they satisfy commutator estimates with gains of regularity uniformly with respect to $N$; see Proposition A.6. These are crucial to parts of the analysis. Notice that because $|1-S(\bm{k})|\leq 1=|1-P_{1/2}(\bm{k})|$ when $\max_{j=1,\ldots,d}|k_{j}|\geq 1/2$ and $1-S(\bm{k})=0=1-P_{1/2}(\bm{k})$ otherwise, we infer from (1.2) the identical spectral convergence estimate
@@
where $A_{j}^{0}=A_{j}(\bm{0})$ and $A_{j}^{1}(\bm{U})=A_{j}(\bm{U})-A_{j}^{0}$. Indeed, applying smooth low-pass filters to linear terms is unnecessary, especially when one uses exponential time integrators (see e.g. Program 27 in [25]). Notice the distinction between (1.5) and (1.6) is only necessary when using the smooth low-pass filter since when using the sharp low-pass filter ${\sf P}_{N}$ one has ${\sf P}_{N}\bm{U}_{N}=\bm{U}_{N}$ and ${\sf P}_{N}$ commutes with $A_{j}^{0}$.
@@ 1.1 Motivation and related works
The aforementioned Boussinesq (resp. Whitham–Boussinesq) systems can be loosely considered as systems generalizing the BBM (resp. Whitham) scalar equation, in the same way the Saint-Venant system discussed in Section 4 generalizes the inviscid Burgers equation. All these equations provide valid approximations of water waves provided (among other assumptions) the shallowness parameter $\mu>0$ satisfies $\mu\ll 1$; see [17, 7].
Results proved in [26] and [6] and adapted to the simplest case of scalar equations are of the form
$\big{\lvert}u_{N}(t,\cdot)\big{\rvert}_{H^{s}}\leq C_{\mu}(t)\big{\lvert}u_{N}\big{|}_{t=0}\big{\rvert}_{H^{s}}\quad\text{ and }\quad\big{\lvert}(u-u_{N})(t,\cdot)\big{\rvert}_{H^{r}}\leq C_{\mu}(t)N^{r-s}$
@@
The rigorous analysis of semi-discretized or fully (space and time) discretized Fourier spectral schemes for semilinear equations is a very rich and active topic, which is impossible to summarize within a few lines; let us simply mention [4, 3] which are particularly relevant as they specifically consider the BBM equation and pay attention to the non-dispersive limit (although together with vanishing nonlinearity, that is the long wave limit). In contrast, to the best of our knowledge, there are only a handful of works dedicated to the rigorous analysis of the Fourier method for quasilinear systems, culminating with the work of Bardos and Tadmor [1] (following [23, 24, 11]; see also the review paper [12]). Here the authors consider the inviscid Burgers equation, as well as the one-dimensional isentropic Euler equation and the incompressible Euler equations. The first two systems belong to the class of equations we study. Specifically, the inviscid Burgers equation (as any scalar equation) has the symmetric structure we employ in Subsection 2.1, while the isentropic Euler equation in Lagrangian coordinates enjoys the Hamiltonian structure discussed in Subsection 3.2. The incompressible Euler equations
@@
does not belong to the class of equations studied here due to the Leray projection operator $\mathbb{P}$ but it would not be difficult to extend our analysis (specifically concerning the symmetric situation) to this system. Let us mention that while our analysis is very similar, the estimates obtained in [1] are not as sharp as the ones obtained in our work due to different choices when performing stability estimates, as they take the form (see Theorem 3.1 therein)
@@
while our results do not feature the second contribution. Quite interestingly, the authors also prove the emergence of spurious oscillations of the semi-discretized solution after the critical time of shock formation for the continuous solution of the inviscid Burgers equation. It would be interesting to study this problem when smooth low-pass filters are used and compare with the spectral viscosity method described therein.
@@ 1.2 Definitions and notations
We use the notation $\left\langle\cdot\right\rangle=(1+\left|\cdot\right|^{2})^{1/2}$ and $\Lambda^{s}=(1-\Delta)^{s/2}$, i.e. the Fourier multiplier with symbol $\left\langle\cdot\right\rangle^{s}$. For real $s\geq 0$, we denote the $L^{2}$-based periodic Sobolev spaces by $H^{s}((2\pi\mathbb{T})^{d})^{n}$:
$H^{s}((2\pi\mathbb{T})^{d})^{n}$ =$=\{\bm{U}\in L^{2}((2\pi\mathbb{T})^{d})^{n},\left|\bm{U}\right|_{H^{s}}<\infty\},\text{ where }$
@@
$\mathcal{T}_{N}\coloneq\operatorname{span}\{\exp({\rm i}\bm{k}\cdot\bm{x}),\bm{k}\in\mathbb{Z}^{d},\left|k_{j}\right|\leq N,j=1,\ldots,d\},\text{ and }\mathcal{T}_{N}^{n}\coloneq\underbrace{\mathcal{T}_{N}\times\ldots\times\mathcal{T}_{N}}_{n\text{ times}},$
@@
We set ${{\sf S}_{N}=\operatorname{Diag}(S_{N}(D))}$ a Fourier multiplier with symbol $S_{N}(\cdot)=S(\cdot/N)$ where $S\in{\mathcal{C}}^{1}(\mathbb{R}^{d})$ is even and satisfies
$\begin{cases}S(\bm{k})=1&\text{if $\max_{j=1,\ldots,d}|k_{j}|\leq 1/2$,}\\ S(\bm{k})=0&\text{if $\min_{j=1,\ldots,d}|k_{j}|\geq 1$,}\\ S(\bm{k})\in{[}0,1{]}&\text{otherwise.}\end{cases}$
and $\left|\cdot\right|\nabla S^{1/2}\in L^{\infty}(\mathbb{R}^{d})$. Apart from these properties, the specific profile of the symbol $S$ is inconsequential.
@@ 2 Symmetric quasilinear systems
Let $s>1+d/2$, and $M>0$. Suppose that for all $j\in\{1,\ldots,d\}$, $A_{j}(\bm{U}_{N})$ satisfies the Assumption A.1 and is self-adjoint. There exists $C>0$ and $T>0$ (depending only on $s$ and $M$) such that for every $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ such that $\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}\leq M$, there exists a unique $\bm{U}\in{\mathcal{C}}(I;H^{s}((2\pi\mathbb{T})^{d})^{n})$ maximal-in-time classical solution to (1.1), and moreover the open time interval $I\supset[0,T/\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}]$ and for all $0\leq t\leq T/\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}$,
@@ 2.1 Discretization with sharp low-pass filters
Recall that the spatial discretization of the problem amounts to finding a solution $\bm{U}_{N}\colon t\mapsto\mathcal{T}_{N}^{n}$ of the problem
@@
We want to show that the semi-discretized solution $\bm{U}_{N}$ to (2.1) converges as $N\to\infty$ towards $\bm{U}$ the corresponding solution to the underlying system (1.1). To do so, we will first show that the semi-discretized solution exists and is bounded on a time interval independent of $N$. This is Proposition 2.2. Then we use this bound to compare the semi-discretized and continuous solution on the interval of existence in Proposition 2.4. Finally, we refine this result by showing that if $N$ is large enough the existence of the semi-discretized solution and the estimate on the difference hold on any compact subset of the interval of existence of the solution to (1.1).
@@
Let $s>d/2+1$ and $M>0$. Suppose that for all $j\in\{1,\ldots,d\}$, $A_{j}(\bm{U}_{N})$ satisfies the Assumption A.1 and is self-adjoint. There exists $C>0$ and $T>0$ depending only on $s$ and $M$ such that for every $N\in\mathbb{N}$ and for every $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ such that $\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}\leq M$, there exists a unique ${\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})}$ maximal-in-time classical solution to (2.1) and $\bm{U}_{N}\big{|}_{t=0}={\sf P}_{N}\bm{U}^{0}$. The open time interval $I_{N}\supset[0,T/\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}]$ and for all $0\leq t\leq T/\big{\lvert}\bm{U}^{0}\big{\rvert}_{H^{s}}$,
@@
Let $s^{\prime}\geq s>d/2+1$. Let $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ be solution to (2.1) on $I_{N}\subset\mathbb{R}$ open time interval. Then $\bm{U}_{N}\in{\mathcal{C}}^{1}(I_{N};H^{s^{\prime}}((2\pi\mathbb{T})^{d})^{n})$ and for any $t\in I_{N}$, one has
@@
Notice $\bm{U}_{N}={\sf P}_{N}\bm{U}_{N}$ and hence we have smoothness in space, $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{\sigma}((2\pi\mathbb{T})^{d})^{n})$ for any $\sigma\in\mathbb{R}$. We infer smoothness in time, $\bm{U}_{N}\in{\mathcal{C}}^{1}(I_{N};H^{\sigma}((2\pi\mathbb{T})^{d})^{n})$, using eq. (2.1), product and composition estimates in $H^{s}((2\pi\mathbb{T})^{d})^{n}$ —Propositions A.3 and A.4— and that ${\sf P}_{N}:H^{s}((2\pi\mathbb{T})^{d})^{n}\to H^{\sigma}((2\pi\mathbb{T})^{d})^{n}$ is bounded. Denote $\dot{\bm{U}}_{N}\coloneq\Lambda^{s^{\prime}}\bm{U}_{N}$. Using that $\bm{U}_{N}$ satisfies the system (2.1), that ${\sf P}_{N}$ is symmetric for the $L^{2}$ inner-product and ${{\sf P}_{N}\dot{\bm{U}}_{N}=\dot{\bm{U}}_{N}}$, and finally- that $A_{j}$ are self-adoint for all $j\in\{1,\ldots,d\}$ and integration by parts, we have
@@
Using the commutator and composition estimates of Propositions A.5,A.4 with $s_{0}=s-1$ for the first contribution, the continuous Sobolev embedding $H^{s}((2\pi\mathbb{T})^{d})^{n}\subset W^{1,\infty}((2\pi\mathbb{T})^{d})^{n}$ (Proposition A.1) for the second as well the Cauchy-Schwarz inequality gives the bound
@@
For each $N>0$, existence and uniqueness of a local-in-time solution $\bm{U}_{N}$ to (2.1) follows from its formulation as a system of ODEs in the Banach space $H^{s}((2\pi\mathbb{T})^{d})^{n}$ (using eq. (2.1), product and composition estimates in $H^{s-1}((2\pi\mathbb{T})^{d})^{n}$ —Propositions A.3 and A.4— and that ${\sf P}_{N}:H^{s-1}((2\pi\mathbb{T})^{d})^{n}\to H^{s}((2\pi\mathbb{T})^{d})^{n}$ is bounded) and the Cauchy–Lipschitz theorem. We denote $I_{N}$ the maximal interval of existence, and notice as in the proof of Lemma 2.3 that $\bm{U}_{N}\in{\mathcal{C}}^{1}(I_{N};H^{s^{\prime}}((2\pi\mathbb{T})^{d})^{n})$ for any $s^{\prime}\geq s$.
@@
Since $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$, $\varphi_{N}\in{\mathcal{C}}(I_{N}\cap\mathbb{R}^{+};\mathbb{R})$ is non-decreasing and $J_{N}=\varphi_{N}^{-1}([0,2\left|\bm{U}^{0}\right|_{H^{s}}])$ is a closed interval. Let us prove that one can set $T>0$ independently of $N$ such that $J_{N}\cap[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$ is an open subset of $[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$. Let $t\in J_{N}\cap[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$ By the second part of the proposition, we can set $C=2C^{*}$ depending only on $s$ and $\left|\bm{U}^{0}\right|_{H^{s}}$ such that
@@
Choosing $T=\ln(3/2)/C$, we find that $\varphi(t)\leq\frac{3}{2}\left|\bm{U}^{0}\right|_{H^{s}}$, and hence —by the continuity of $\varphi_{N}$— there exists $\delta>0$ such that $[t-\delta,t+\delta]\in J_{N}$. This concludes the proof that $J_{N}\cap[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$ is a non-empty, connected, closed and open subset of $[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$. Hence $I_{N}\supset J_{N}\supset[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$ and estimate (2.3) concludes the proof. ∎

### v2

@@ 1 Introduction
Assuming that the system (1.1) has only polynomial nonlinearities may seem an over-restrictive assumption. This assumption is motivated by two considerations. Firstly, the Fourier spectral method may be efficiently implemented only within this framework; see Remark 1.2 below. Secondly, as far as we know, the literature lacks a theory analogous to para-differential calculus (see e.g.  [19, Chapter 5]) on Sobolev spaces of periodic functions, which would provide the composition estimates in Proposition A.4 for general composition functions. All our results apply without assuming polynomial nonlinearities if Proposition A.4 holds without that assumption.
We shall assume further additional structural assumptions on the system (1.1), depending on the needs. In particular we shall always consider Friedrichs-symmetrizable systems, which guarantees the hyperbolicity of the system, and local-in-time well-posedness of the initial-value problem in $L^{2}$-based Sobolev spaces of sufficiently high regularity index, $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ with $s>d/2+1$ (see Subsection 1.2). For simplicity, we consider in this work $2\pi$-periodic functions in all spatial directions. The results extend straightforwardly to more general periodic frameworks, and analogous results in the $n$-dimensional Euclidean space could be obtained with some simple adaptations.
@@
Let ${\sf P}_{N}\colon L^{2}((2\pi\mathbb{T})^{d})^{n}\to\mathcal{T}_{N}^{n}$ be the $L^{2}$-projection operator onto $\mathcal{T}_{N}^{n}$: ${\sf P}_{N}=\operatorname{Diag}(P_{N}(D))$, where ${\sf P}_{N}$ is a Fourier multiplier with symbol $P_{N}(\cdot)=\bm{1}_{\llbracket-N,N\rrbracket^{d}}(\cdot)$. We refer to ${\sf P}_{N}$ as a sharp low-pass filter. For any $0\leq r\leq s$ and $\bm{U}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ Sobolev space of order $s$, ${\sf P}_{N}$ satisfies (see e.g. [5, (5.1.10),(5.8.4)])
@@
We do not discuss in this work the full time-space discretization of (1.1), that is well-suited numerical time integrators for (1.3). The implementation we used for numerical experiments is described in more detail in Subsections 4.2 and 4.3.
@@
Our results will depend on the structure of the system. As we shall see, the solution to (1.3) converges towards the corresponding solution to the problem (1.1) (assuming sufficient regularity) whenever the system is symmetric. If the system is only symmetrizable, the situation is more complicated. In order to deal with this situation we consider smooth low-pass filters, ${\sf S}_{N}\colon L^{2}((2\pi\mathbb{T})^{d})^{n}\to\mathcal{T}_{N}^{n}$ where ${\sf S}_{N}=\operatorname{Diag}(S_{N}(D))$} is a Fourier multiplier with symbol $S_{N}(\cdot)=S(\cdot/N)$ where $S$ is even and satisfies
$\begin{cases}S(\bm{k})=1&\text{if $\max_{j=1,\ldots,d}|k_{j}|\leq 1/2$,}\\ S(\bm{k})=0&\text{if $\min_{j=1,\ldots,d}|k_{j}|\geq 1$,}\\ S(\bm{k})\in{[}0,1{]}&\text{otherwise,}\end{cases}$
and $S^{1/2}$ is Lipschitz-continuous. When $d=1$, an example of such a function is $S_{1}(\cdot)\coloneq\max\big(0,\min\big(1,2-2|\cdot|\big)\big)^{2}$. When $d\geq 2$, one can set $S_{d}((k_{1},\dots,k_{d}))\coloneq S_{1}(k_{1})\times\cdots\times S_{1}(k_{d})$. The advantage of such smooth low-pass filters is that —contrarily to the sharp low-pass filter— they satisfy commutator estimates with gains of regularity uniformly with respect to $N$; see Proposition A.6. These are crucial to parts of the analysis. Notice that because $|1-S(\bm{k})|\leq 1=|1-P_{1/2}(\bm{k})|$ when $\max_{j=1,\ldots,d}|k_{j}|\geq 1/2$ and $1-S(\bm{k})=0=1-P_{1/2}(\bm{k})$ otherwise, we infer from (1.2) the corresponding spectral convergence estimate
@@
where $A_{j}^{0}=A_{j}(\bm{0})$ and $A_{j}^{1}(\bm{U})=A_{j}(\bm{U})-A_{j}^{0}$. Indeed, applying smooth low-pass filters to linear terms is unnecessary, especially when one uses exponential time integrators (see e.g. Program 27 in [25]). Notice the distinction between (1.5) and (1.6) is only necessary when using smooth low-pass filters since when using the sharp low-pass filter ${\sf P}_{N}$ one has ${\sf P}_{N}\bm{U}_{N}=\bm{U}_{N}$ and ${\sf P}_{N}$ commutes with $A_{j}^{0}$.
@@ 1.1 Motivation and related works
The aforementioned Boussinesq (resp. Whitham–Boussinesq) systems can be loosely considered as systems extending the BBM (resp. Whitham) scalar equation, in the same way the Saint-Venant system discussed in Section 4 extends the inviscid Burgers equation. All these equations provide valid approximations of water waves provided (among other assumptions) the shallowness parameter $\mu>0$ satisfies $\mu\ll 1$; see [17, 7].
Results proved in [26] and [6] and adapted to the simplest case of scalar equations take the form
$\big\lvert u_{N}(t,\cdot)\big\rvert_{H^{s}}\leq C_{\mu}(t)\big\lvert u_{N}\big|_{t=0}\big\rvert_{H^{s}}\quad\text{ and }\quad\big\lvert(u-u_{N})(t,\cdot)\big\rvert_{H^{r}}\leq C_{\mu}(t)N^{r-s}\qquad\text{ for any $s>1/2$ and $0\leq r\leq s$,}$
@@
The rigorous analysis of semi-discretized or fully (space and time) discretized Fourier spectral schemes for semilinear equations is a very rich and active topic, which is impossible to summarize within a few lines; let us simply mention [4, 3] which are particularly relevant as they specifically consider the BBM equation and pay attention to the non-dispersive limit (although together with vanishing nonlinearity, that is the long wave limit). In contrast, to the best of our knowledge, there are only a handful of works dedicated to the rigorous analysis of the Fourier method for quasilinear systems, culminating with the work of Bardos and Tadmor [1] (following [23, 24, 11]; see also the review paper [12]). Here the authors consider the inviscid Burgers equation, as well as the one-dimensional isentropic Euler equation and the incompressible Euler equations. The first two systems belong to the class of equations we study. Specifically, the inviscid Burgers equation has the symmetric structure we employ in Subsection 2.1, while the isentropic Euler equation in Lagrangian coordinates enjoys the Hamiltonian structure discussed in Subsection 3.2. The incompressible Euler equations
@@
does not belong to the class of equations studied in this work due to the presence of the Leray projection operator $\mathbb{P}$ but it would not be difficult to extend our analysis (specifically concerning the symmetric situation) to this system. Let us mention that while our analysis is very similar, the estimates obtained in [1] are not as sharp as the ones obtained in our work due to different choices when performing stability estimates, as the former take the form (see Theorem 3.1 therein)
@@
while our results do not feature the second contribution. Quite interestingly, the authors in [1] also prove the emergence of spurious oscillations of the semi-discretized solution after the critical time of shock formation for the continuous solution of the inviscid Burgers equation. It would be interesting to study this problem when smooth low-pass filters are used and compare with the spectral viscosity method described therein.
@@ 1.2 Definitions and notations
We use the notation $\left\langle\cdot\right\rangle=(1+\left|\cdot\right|^{2})^{1/2}$ and $\Lambda^{s}=(1-\Delta)^{s/2}$, i.e. the Fourier multiplier with symbol $\left\langle\cdot\right\rangle^{s}$ (see e.g. [19] for Fourier multipliers). For real $s\geq 0$, we denote the $L^{2}$-based periodic Sobolev spaces by $H^{s}((2\pi\mathbb{T})^{d})^{n}$:
$H^{s}((2\pi\mathbb{T})^{d})^{n}$ $=\{\bm{U}\in L^{2}((2\pi\mathbb{T})^{d})^{n},\left|\bm{U}\right|_{H^{s}}<\infty\},\text{ where }$
@@
$\mathcal{T}_{N}\coloneq\operatorname{span}\{\exp({\rm i}\bm{k}\cdot\bm{x}),\bm{k}=(k_{1},\dots,k_{d})\in\mathbb{Z}^{d},\left|k_{j}\right|\leq N,j=1,\ldots,d\},\text{ and }\mathcal{T}_{N}^{n}\coloneq\underbrace{\mathcal{T}_{N}\times\ldots\times\mathcal{T}_{N}}_{n\text{ times}},$
@@
We set ${\sf S}_{N}=\operatorname{Diag}(S_{N}(D))$} a Fourier multiplier with symbol $S_{N}(\cdot)=S(\cdot/N)$ where $S$ is even and satisfies
$\begin{cases}S(\bm{k})=1&\text{if $\max_{j=1,\ldots,d}|k_{j}|\leq 1/2$,}\\ S(\bm{k})=0&\text{if $\min_{j=1,\ldots,d}|k_{j}|\geq 1$,}\\ S(\bm{k})\in{[}0,1{]}&\text{otherwise,}\end{cases}$
and $S^{1/2}$ is Lipschitz-continuous. Apart from these properties, the specific profile of the symbol $S$ is inconsequential.
@@ 2 Symmetric quasilinear systems
Let $s>1+d/2$, and $M>0$. Suppose that for all $j\in\{1,\ldots,d\}$, $A_{j}(\cdot)$ satisfies the Assumption A.1 and is self-adjoint. There exists $C>0$ and $T>0$ (depending only on $s$ and $M$) such that for every $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ such that $\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}\leq M$, there exists a unique $\bm{U}\in{\mathcal{C}}(I;H^{s}((2\pi\mathbb{T})^{d})^{n})$ maximal-in-time classical solution to (1.1), and moreover the open time interval $I\supset[0,T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}]$ and for all $0\leq t\leq T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}$,
@@ 2.1 Discretization with sharp low-pass filters
Recall that the spatial discretization of the problem (1.1) amounts to finding a solution $\bm{U}_{N}\colon t\mapsto\mathcal{T}_{N}^{n}$ to the problem
@@
We want to show that the semi-discretized solutions $\bm{U}_{N}$ to (2.1) converge as $N\to\infty$ towards $\bm{U}$ the corresponding solution to the underlying system (1.1). To do so, we will first show that the semi-discretized solutions exist and are bounded on a time interval independent of $N$. This is Proposition 2.2. Then we use this bound to compare the semi-discretized and continuous solution on the interval of existence in Proposition 2.4. Finally, we refine this result by showing that if $N$ is large enough the existence of the semi-discretized solution and the estimate on the difference hold on any compact subset of the interval of existence of the solution to (1.1).
@@
Let $s>d/2+1$ and $M>0$. Suppose that for all $j\in\{1,\ldots,d\}$, $A_{j}(\cdot)$ satisfies the Assumption A.1 and is self-adjoint. There exists $C>0$ and $T>0$ depending only on $s$ and $M$ such that for every $N\in\mathbb{N}$ and for every $\bm{U}^{0}\in H^{s}((2\pi\mathbb{T})^{d})^{n}$ such that $\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}\leq M$, there exists a unique $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$} maximal-in-time classical solution to (2.1) and $\bm{U}_{N}\big|_{t=0}={\sf P}_{N}\bm{U}^{0}$. The open time interval $I_{N}\supset[0,T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}]$ and for all $0\leq t\leq T/\big\lvert\bm{U}^{0}\big\rvert_{H^{s}}$,
@@
Let $s^{\prime}\geq s>d/2+1$. Suppose that for all $j\in\{1,\ldots,d\}$, $A_{j}(\cdot)$ satisfies the Assumption A.1 and is self-adjoint. Let $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$ be solution to (2.1) on $I_{N}\subset\mathbb{R}$ an open time interval. Then $\bm{U}_{N}\in{\mathcal{C}}^{1}(I_{N};H^{s^{\prime}}((2\pi\mathbb{T})^{d})^{n})$ and for any $t\in I_{N}$, one has
@@
Notice $\bm{U}_{N}={\sf P}_{N}\bm{U}_{N}$ and hence we have smoothness in space, $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{\sigma}((2\pi\mathbb{T})^{d})^{n})$ for any $\sigma\in\mathbb{R}$. We infer smoothness in time, $\bm{U}_{N}\in{\mathcal{C}}^{1}(I_{N};H^{\sigma}((2\pi\mathbb{T})^{d})^{n})$, using eq. (2.1), product and composition estimates in $H^{s}((2\pi\mathbb{T})^{d})^{n}$ —Propositions A.3 and A.4— and that ${\sf P}_{N}:H^{s}((2\pi\mathbb{T})^{d})^{n}\to H^{\sigma}((2\pi\mathbb{T})^{d})^{n}$ is bounded. Denote $\dot{\bm{U}}_{N}\coloneq\Lambda^{s^{\prime}}\bm{U}_{N}$. Using that $\bm{U}_{N}$ satisfies the system (2.1), that ${\sf P}_{N}$ is symmetric for the $L^{2}$ inner-product and ${\sf P}_{N}\dot{\bm{U}}_{N}=\dot{\bm{U}}_{N}$}, and finally that $A_{j}$ are self-adoint for all $j\in\{1,\ldots,d\}$ and integration by parts, we have
@@
Using the commutator and composition estimates of Propositions A.5,A.4 with $s_{0}=s-1$ for the first contribution, the continuous Sobolev embedding $H^{s}((2\pi\mathbb{T})^{d})^{n}\subset W^{1,\infty}((2\pi\mathbb{T})^{d})^{n}$ (Proposition A.1) for the second as well the Cauchy–Schwarz inequality gives the bound
@@
For each $N>0$, existence and uniqueness of a local-in-time solution $\bm{U}_{N}$ to (2.1) follows from its formulation as a system of ODEs in the Banach space $H^{s}((2\pi\mathbb{T})^{d})^{n}$ (using eq. (2.1), product and composition estimates in $H^{s-1}((2\pi\mathbb{T})^{d})^{n}$ —Propositions A.3 and A.4— and that ${\sf P}_{N}:H^{s-1}((2\pi\mathbb{T})^{d})^{n}\to H^{s}((2\pi\mathbb{T})^{d})^{n}$ is bounded) and the Picard–Lindelöf theorem. We denote $I_{N}$ the maximal interval of existence, and notice as in the proof of Lemma 2.3 that $\bm{U}_{N}\in{\mathcal{C}}^{1}(I_{N};H^{s^{\prime}}((2\pi\mathbb{T})^{d})^{n})$ for any $s^{\prime}\geq s$.
@@
Since $\bm{U}_{N}\in{\mathcal{C}}(I_{N};H^{s}((2\pi\mathbb{T})^{d})^{n})$, $\varphi_{N}\in{\mathcal{C}}(I_{N}\cap\mathbb{R}^{+};\mathbb{R})$ is non-decreasing and $J_{N}=\varphi_{N}^{-1}([0,2\left|\bm{U}^{0}\right|_{H^{s}}])$ is a closed interval. Let us prove that one can set $T>0$ independently of $N$ such that $J_{N}\cap[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$ is a non-empty open subset of $0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$. Notice $0\in J_{N}$. Let $t\in J_{N}\cap[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$ . By the second part of the proposition, we can set $C=2C^{*}$ depending only on $s$ and $\left|\bm{U}^{0}\right|_{H^{s}}$ such that
@@
Choosing $T=\ln(3/2)/C$, we find that $\varphi_{N}(t)\leq\frac{3}{2}\left|\bm{U}^{0}\right|_{H^{s}}$, and hence —by the continuity and monotonicity of $\varphi_{N}$— there exists $\delta>0$ such that $0,t+\delta]\in J_{N}$. This concludes the proof that $J_{N}\cap[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$ is a non-empty, connected, closed and open subset of $0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$. Hence $I_{N}\supset J_{N}\supset[0,T/\left|\bm{U}^{0}\right|_{H^{s}}]$ and estimate (2.3) concludes the proof. ∎